OpenStudy (anonymous):

Select the expression below that is identical to this equation: (I'll post it with the equation maker thingy :3)

OpenStudy (anonymous):

\[\tan (-\theta)\tan (\frac{ \pi }{ 2 }-\theta)\]

OpenStudy (anonymous):

The options to chose from are as follows: A.) 1 B.) \[\cot ^{2}\theta \] C.)\[\tan ^{2}\theta \] D.)0 E.) -1

OpenStudy (anonymous):

I don't understand how to simplify the original equation enough to get an answer. I know pi/2 = 90 degrees, but I'm not sure how that fits into the whole scheme of things.

OpenStudy (anonymous):

okay so one of my identities is tan(pi/2 -x)=cot x so now i have \[\tan (-\theta) \cot \theta\]

OpenStudy (anonymous):

cot is just the reciprocal of tan and since tan has a negative theta and cot a positive theta then they'll all cancel out cause they equal each other. equaling 1! :D

OpenStudy (anonymous):

no wait, because tan(- theta) equals cot theta it should be negative one right?